OpenStudy (anonymous):

Every month Jon saves $350.00. If he earns 2.5% interest on this amount, how much will he have saved in a year? A. $1,050.00 B. $2,795.00 C. $4,200.00 D. $4,305.00

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (jdoe0001):

I'm assuming they're saying, that in every month he earns 2.5% so he really deposits in the bank $350 PLUS 2.5% of $350 that is 350 + 350(0.025) every month so in a year, well there are 12 months in a year, so 12 times that much

OpenStudy (anonymous):

okay I got D!

OpenStudy (jdoe0001):

yeah, I got the same :)

OpenStudy (wolf1728):

I think it might be referring to a monthly annuity. The formula for which is attached,

OpenStudy (anonymous):

im confused now.. so are you saying you may think jdoes anser is wrong?

OpenStudy (wolf1728):

I'm just saying maybe. I'm not sure. It's just that the problem seems as if it is referring to an annuity.

OpenStudy (anonymous):

@amistre64 would you say D is correct?

OpenStudy (amistre64):

it is an annuity

OpenStudy (amistre64):

and unless stated otherwise, the interest rate is given as an annual rate and would have to be modified by /12 ... for 12 periods

OpenStudy (amistre64):

350(1+.025/12)^12 + 350(1-(1+.025)^(12))/(1-(1.025/12))

OpenStudy (amistre64):

... one of my typos abounds im sure

OpenStudy (jdoe0001):

hehe

OpenStudy (anonymous):

im so confused here lol

OpenStudy (amistre64):

350(1+.025/12)^12 + 350(1-(1+.025/12)^(12))/(1-(1+.025/12)) thats better :)

OpenStudy (amistre64):

I get 4607, and since the rate is so low to start with ... im assuming D is correct regardless

OpenStudy (anonymous):

4607.31173314

OpenStudy (wolf1728):

Now that I notice, the multiple choice answers are even amounts and an annuity is not something that would calculate to an even amount.

OpenStudy (amistre64):

multiple choice options tend to round :)

OpenStudy (wolf1728):

Okay, as I said it sure sounds like an annuity to me.

OpenStudy (amistre64):

Mo = 350 M1 = 350(1+.025/12) + 350 M2 = 350(1+.025/12)^2 + 350(1+.025/12) + 350 M3 = 350(1+.025/12)^3 +350(1+.025/12)^2 + 350(1+.025/12) + 350 letting k clean this up we get a geometric sum Mn = 350 + 350k + 350k^2 + ... + 350k^n

OpenStudy (amistre64):

1+k+...+k^n -k-...-k^n - k^(n+1) -------------------- [1-k^(n+1)]/(1-k)

OpenStudy (wolf1728):

I get 4,257.31 from a calculator located here: http://www.coolmath.com/calculators/calculator-annuity-1.html

OpenStudy (amistre64):

350(1-k^(13))/(1-k), k=1+.025/12 http://www.wolframalpha.com/input/?i=350%281-k%5E%2813%29%29%2F%281-k%29%2C+k%3D1%2B.025%2F12 4607 still :)

OpenStudy (amistre64):

even at best ... 350 for 12 months is 4200, so with interest it has to be greater than that

OpenStudy (wolf1728):

Well Kaitlin, I'm sorry about bringing up the concept of an annuity but it's something to consider if your teacher mentions it.

OpenStudy (anonymous):

dont be confused. you will receive different answers for compounding using daily, weekly, monthly, semiannual, and annually. it all depends on how it is compounded. you may want to check with your teacher. the general rule when none is given is to compound annually.